QUESTION 14: Which of the following ratios would be most useful to a creditor to determine an ongoing concern’s (ongoing company’s) creditworthiness?

current ratio

quick ratio

total asset turnover

times interest earned

QUESTION 21: Trico Windshield Wipers Corporation has a 24% return on equity. The debt ratio is 35%. If the total asset turnover is 1.5X, what is the firm's profit margin?

13.85%

11.31%

14.25%

10.40%

none of these
Answer #1

14.

times interest earned

this is related to meeting the debt obligations of the company, thus related to creditworthiness

15.

debt ratio=35%

(debt/assets)=35%

1-(debt/assets)=1-35%

equity/assets=65%

equity multiplier=1/65%=1.5385

what is the firm's profit margin

=return on equity/(asset turnover*equity multiplier)

=24%/(1.5*1.5385)

=10.40%